1. The activation energy for the gas phase decomposition of chloroethane is 254 kJ.

CH3CH2Cl--->C2H4 + HCl
The rate constant at 750 K is 7.99×10-4 /s. The rate constant will be 5.61×10-3 /s at_____ K

2.

The activation energy for the gas phase decomposition of chloroethane is 254 kJ.

CH3CH2Cl--->C2H4 + HCl
The rate constant at 756 K is 1.09×10-3 /s. The rate constant will be___ /s at 786 K.
Answer #1

1.    Ea   = 254KJ   = 254000J

       K1   = 7.99*10^-4 /S

       T1    = 750K

         K2   = 5.61*10^-3 /S

          T2 =

log(K2/K1)   = Ea/2.303*R [ 1/T1 - 1/T2]

log(5.61*10^-3/7.99*10^-4) = 254000/2.303*8.314 [ 1/750 - 1/T2]

0.846          = 13265.6[ 1/750 - 1/T2]

[ 1/750 - 1/T2]    = 0.846/13265.6

[ 1/750 - 1/T2]   = 6.37*10^-5

[ 0.00133 - 1/T2]   = 0.0000637

-1/T2                    = 0.0000637-0.00133

-1/T2                    = -0.0012663

T2                         = 1/0.0012663

   T2                        = 789.7K

2.

   Ea   = 254KJ   = 254000J

       K1   = 1.09*10^-3 /S

       T1    = 756K

         K2   =

          T2 =786K

log(K2/K1)   = Ea/2.303*R [ 1/T1 - 1/T2]

log(K2/1.09*10^-3) = 254000/2.303*8.314 [ 1/756 - 1/786]

  log(K2/1.09*10^-3)       = 13265.6[ 1/756 - 1/786]

log(K2/1.09*10^-3)      = 13265.6(0.00132-0.00127)

log(K2/1.09*10^-3)      = 0.66328

K2/1.09*10^-3           = 10^0.66328

K2/1.09*10^-3           = 4.6055

    K2                         = 1.09*10^-3*4.6055    = 0.005

K2   = 5*10^-3/S >>>>>answer